OpenStudy (anonymous):

Find the length of the unknown side (hypotenuse). Round to three decimal places if necessary. Leg = 2 m, leg = 8m

OpenStudy (anonymous):

Show steps please and thank you! :)

OpenStudy (anonymous):

use pythagorean theorem a^2 + b^2 = C^2 2^2 + 8^2 = c^2 4+64 = c^2 68 = c^2 Then sqrt room each side c = 2 times the sqrt of 17

OpenStudy (anonymous):

Where did 17 come from??

OpenStudy (anonymous):

well, i did factor trees, and if you start at teh top, its 68, which breaks into 2* 34, which breaks into 2*17....you keeep two, and 17..which makes two times the sqrt of 17

OpenStudy (anonymous):

I am so confused. Lol

OpenStudy (anonymous):

do you understand up to the part that says 68 = c^2?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok, ill draw it

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):